# WorldMem: Long-term Consistent World Simulation with Memory

### v1

@@ Abstract
World simulation has gained increasing popularity due to its ability to model virtual environments and predict the consequences of actions. However, the limited temporal context window often leads to failures in maintaining long-term consistency, particularly in preserving 3D spatial consistency. In this work, we present WorldMem, a framework that enhances scene generation with a memory bank consisting of memory units that store memory frames and states (e.g., poses and timestamps). By employing a memory attention mechanism that effectively extracts relevant information from these memory frames based on their states, our method is capable of accurately reconstructing previously observed scenes, even under significant viewpoint or temporal gaps. Furthermore, by incorporating timestamps into the states, our framework not only models a static world but also captures its dynamic evolution over time, enabling both perception and interaction within the simulated world. Extensive experiments in both virtual and real scenarios validate the effectiveness of our approach.
Figure 1: WorldMem enables long-term consistent world simulation with an integrated memory mechanism. (a) Previous world simulation methods typically face the problem of inconsistent world due to limited temporal context window size. (b) WorldMem  empowers the agent to explore diverse and consistent worlds with an expansive action space, e.g., crafting environments by placing objects like pumpkin light or freely roaming around. Most importantly, after exploring for a while and glancing back, we find the objects we placed are still there, with the inspiring sight of the light melting surrounding snow, testifying to the passage of time. Red and green boxes indicate scenes that should be consistent. Project page at https://xizaoqu.github.io/worldmem.
@@ 1 Introduction
World simulation has gained significant attention for its ability to model environments and predict the outcomes of actions [2, 5, 1, 7, 27, 34]. Recent advances in video diffusion models have further propelled this field, enabling high-fidelity rollouts of potential future scenarios based on user actions, such as navigating through an environment or interacting with objects. These capabilities make world simulators particularly promising for applications in autonomous navigation [7, 2] and as an alternative to traditional game engines [5, 27].
Despite impressive progress, existing methods suffer from a key limitation: they struggle to maintain long-term consistency. A prominent issue is the lack of 3D spatial consistency: As illustrated in Figure 1(a), when the viewpoint moves in one direction and then returns, the content of the environment has changed. This inconsistency arises because video diffusion models operate within limited temporal windows, restricting their ability to retain information over extended time periods. Ideally, the model should be able to recall the generated environment and past events even across long temporal gaps. However, simply increasing the temporal window is impractical due to the prohibitive memory and computational costs.
To address this limitation, we introduce WorldMem, a novel framework that enhances long-term consistency in video-based world simulators through a dedicated memory mechanism. Our core idea is to continuously store visual and state information in an external memory bank and retrieve the most relevant memories for generating new frames. By removing the constraint of a bounded time horizon, this memory mechanism allows the model to capture and reuse historical cues, ensuring better preservation of environment details and past events.
WorldMem is built upon a Conditional Diffusion Transformer (CDiT) [28] that integrates external action signals to guide first-person viewpoint generation. By leveraging Diffusion Forcing (DF) [3] during training, the model supports autoregressive generation, enabling extended temporal simulation. Unlike previous methods that store past content either as explicit 3D reconstructions [24, 36] or within implicit network weights [1, 17], we instead maintain a memory bank composed of memory units – each containing a past frame and its corresponding state. This approach offers both flexibility and precision: We avoid rigid 3D reconstructions while retaining the fine-grained visual details necessary for accurate scene synthesis. The memory bank is continuously updated throughout generation, and a dedicated retrieval strategy effectively retrieves relevant units to guide future generation steps.
To effectively incorporate the retrieved memory units, we adopt the flexible DF paradigm, treating memory frames as “clear” latents alongside noisy latents in the denoising loop. This design obviates the need for additional conditioning techniques [15, 48, 43] and prevents information loss that may occur when passing through separate conditioning modules. Following [8, 38], we model information exchange between memory frames and the frames currently being generated through attention [35], where the current frames act as queries and the memory frames serve as keys and values. To ensure robust information retrieval, we represent states with specialized embeddings that enrich both queries and keys. Specifically, we employ Plücker embeddings [31] for dense pose representation and incorporate relative embeddings to facilitate efficient learning.
@@
We evaluate WorldMem on a customized Minecraft benchmark [6] and on RealEstate10K [51]. The Minecraft benchmark includes diverse terrains (e.g., plains, savannas, and deserts) and various action modalities (movement, viewpoint control, and event triggers). Extensive experiments show that WorldMem significantly improves 3D spatial consistency, enabling robust viewpoint reasoning and high-fidelity scene reconstructions, as shown in Figure 1(b). Furthermore, in dynamic environments, WorldMem accurately tracks and follows evolving events and environment changes, demonstrating its ability to both perceive and interact with the generated world.
Overall, WorldMem’s enhanced capacity to maintain long-term consistency represents a critical advancement for interactive world simulators, paving the way for more accurate, persistent, and immersive virtual environments. We hope our promising results will inspire future research on memory-based consistent world simulation.
@@ 2 Related Work
Video diffusion model. With the rapid advancement of diffusion models [33, 28, 3], video generation has made significant strides [37, 39, 4, 9, 26, 20, 44]. The field has evolved from traditional U-Net-based architectures [37, 4, 9] to Transformer-based frameworks [26, 25, 50], enabling video diffusion models to generate highly realistic and temporally coherent videos. Recently, autoregressive video generation [3, 21, 15] has emerged as a promising approach to extend video length, theoretically indefinitely. Notably, Diffusion Forcing [3] introduces a per-frame noise-level denoising paradigm. Unlike the full-sequence paradigm, which applies a uniform noise level across all frames, per-frame noise-level denoising offers a more flexible approach, enabling autoregressive generation.
@@
Interactive world simulation. World simulation aims to model an environment by predicting the next state given the current state and action. This concept has been extensively explored in the construction of world models [11] for agent learning [10, 12, 13, 18]. With advances in video generation, high-quality world simulation with robust control has become feasible, leading to numerous works focusing on interactive world simulation [2, 5, 1, 7, 27, 34, 47]. These approaches enable agents to navigate generated environments and interact with them based on external commands.
@@
However, due to context window limitations, such methods tend to forget previously generated content, leading to inconsistencies in the simulated world, particularly in maintaining 3D spatial coherence.
Consistent world simulation. Ensuring the consistency of a generated world is crucial for effective world simulation. Existing approaches can be broadly categorized into two types: explicit reconstructions and implicit conditioning.
One mainstream approach to ensuring consistency is to explicitly reconstruct the generated world into a 3D representation [24, 8, 36, 29, 46, 45, 23]. While this strategy can reliably maintain consistency, it imposes strict constraints on flexibility: Once the world is reconstructed, modifying or interacting with it becomes challenging.
@@
Another research direction focuses on implicit learning. Methods like [1, 34] ensure consistency by overfitting to predefined scenarios (e.g., specific CS:GO or DOOM maps), limiting scalability. StreamingT2V [15] maintains long-term consistency by continuing on both global and local visual contexts from previous frames, while SlowFastGen [17] progressively trains LoRA [19] modules for memory recall. However, these methods rely on abstract representations, making accurate scene reconstruction challenging. In contrast, our approach retrieves information from previously generated frames and their states, ensuring world consistency without overfitting to specific scenarios.
@@ 3 WorldMem
This section details the methodology of WorldMem. Sec. 3.1 introduces the relevant preliminaries, while Sec. 3.2 describes the interactive world simulator serving as our baseline. Sec. 3.3 presents the core of our proposed memory mechanism, followed by Sec. 3.4 covering the memory retrieval strategies and Sec. 3.5 discussing the key designs for state embeddings in memory blocks.
@@ 3.1 Preliminary
Autoregressive video generation. Autoregressive video generation aims to extend videos over the long term by predicting frames sequentially [22, 41]. While various methods exist for autoregressive generation, Diffusion Forcing (DF) [3] provides a neat and effective approach to achieve this. Specifically, DF introduces per-frame noise levels $k_{t}$:
@@ 3.2 Interactive World Simulation
Before introducing the memory mechanism, we first outline the construction of a world simulator that incorporates actions as additional inputs within a conditional diffusion transformer framework. Following previous works [5], we adopt a conditional diffusion transformer as the baseline for interactive world simulation. Specifically, we adopt a DiT [28] architecture for video generation and use DF [3] for autoregressive generation. As illustrated in Figure 2 (a), our model comprises multiple DiT blocks with spatial and temporal modules for spatiotemporal reasoning. In the temporal module, causal attention ensures each frame only attends to preceding frames.
To further support interactive world simulation, we integrate interactive modules that incorporate external control signals into the models. The primary control signal is action [27, 5, 47], which is crucial in guiding the simulation. In our Minecraft experiments, the action space consists of $25$ dimensions, including movements (left, right, forward, back), view controls (look up, look down), and event triggers (object placement, item switching). To condition the model on these actions, we employ a multi-layer perceptron (MLP) to project the control signals into the embedding space. As shown in Figure 2 (c), we add the action embeddings and denoising timestep embedding together and inject the information to temporal blocks by AdaLN  [42], following the paradigm of previous approaches [2, 5]. Please note that we also inject timestep embeddings to spatial blocks in the same way, but we do not highlight that in Figure 2 for simplicity. Likewise, default attention settings like residual connections, multi-head attention, and feedforward networks are also omitted in Figure 2 for simplicity.
@@
With the combination of conditional DiT [28] and DF [3] as a baseline, we can effectively simulate a long-term interactive world. However, due to the high computational cost of video generation, the context window remains limited. Consequently, content outside this window is forgotten, leading to inconsistencies in long-term generation [5].
3.3 Condition on Memory
We introduce a memory mechanism to prevent the model from forgetting content outside the context window. Intuitively, this mechanism functions as a persistent external buffer, allowing the model to revisit past frames and maintain a continuous record of the evolving scene. Specifically, we design a memory bank that continuously stores previously generated results. The memory is structured as a series of memory units $\mathbf{M}=\{\mathbf{I},\mathbf{p},\mathbf{t}\}$, which include visual content in the form of memory frames $\mathbf{I}$, along with state information such as poses $\mathbf{p}$ and timestamps $\mathbf{t}$.
However, effectively utilizing these memory units to guide generation is non-trivial. Unlike existing conditioning methods [15, 17], which condition on high-level and abstract features, our objective is to accurately reconstruct previously seen scenarios, even under significant viewpoint changes or scene changes. This requires the conditioning module to reason about the relationships between past results and the current generation.
To achieve this, we adopt the cross-attention mechanism [35] to guide information exchange between the denoising frames and the memory frames. Let $\mathbf{X}\in\mathbb{R}^{l_{1}\times d}$ be the denoising frames (queries) and $\mathbf{M}\in\mathbb{R}^{l_{2}\times d}$ be the memory frames (keys and values), where $l_{1}$ and $l_{2}$ are the flattened lengths of $\mathbf{X}$ and $\mathbf{M}$, respectively, and $d$ is the channel dimension. We first incorporate the state embeddings $\mathbf{E}_{X}$ and $\mathbf{E}_{M}$ by summation:
$\mathbf{\tilde{X}}=\mathbf{X}+\mathbf{E}_{X},\quad\mathbf{\tilde{M}}=\mathbf{M}+\mathbf{E}_{M}.$ (3)
We then compute cross-attention as:
$\mathbf{X}^{\prime}=\text{CrossAttn}(\text{Q}=p_{q}(\mathbf{\tilde{X}}),\text{K}=p_{k}(\mathbf{\tilde{M}}),\text{V}=p_{v}(\mathbf{M})),$ (4)
where $\mathbf{X}^{\prime}$ is the output after attending to relevant information in the memory frames, $p_{q}$, $p_{k}$, and $p_{v}$ are attention projectors. This allows the network to focus on memory content that is most pertinent to the current generation step. We will further elaborate on how $\mathbf{E}_{X}$ and $\mathbf{E}_{M}$ are computed in Sec. 3.5.
To incorporate memory units into different DiT blocks, we leverage the advantages of DF by treating memory frames as clear frames and feeding them into the diffusion pipeline alongside input noisy frames, as shown in Figure 2 (b). For the training process, all frames in the temporal context window have different levels of noise, and these memory frames have a clear level of noise (indicated as $k_{min}$):
$\mathbf{x}_{i}^{k},k\begin{cases}\sim\mathcal{P}(\mathbf{k}),&i\in\mathbf{TF}\\ =k_{\text{min}},&i\in\mathbf{MF}\end{cases}$ (5)
where $\mathbf{x}_{i}^{k}$ indicates the $i^{\text{th}}$ frame with noise level $k$, $\mathbf{k}$ is the set of possible noise levels, $\mathbf{TF}$ is the set of frame indices within temporal context window, $\mathbf{MF}$ is the set of frame indices in the memory frame part. For inference,
$\mathbf{x}_{i}^{k},k=\begin{cases}k_{\text{max}},&i\in\mathbf{GF}\\ k_{\text{min}},&\text{otherwise}\end{cases}$ (6)
where $\mathbf{GF}$ is the set of indices of generated frames, $k_{\text{max}}$ indicates the max level of noise.
We ensure that these memory frames influence only the memory block by masking them in the temporal block. As shown in Figure 2 (c), given the memory frame length as $L_{M}$, the mask can be noted as:
$A_{\text{mask}}(i,j)=\begin{cases}1,&\text{if }i\leq L_{M}\text{ and }j=i\\ 1,&\text{if }i>L_{M}\text{ and }j\leq i\\ 0,&\text{otherwise}\end{cases}$ (7)
Within the memory block, we use cross-attention to retrieve information from memory frames and guide the generation of the current frame.
3.4 Memory Retrieve
Since the number of memory frames available for conditioning is limited, an efficient strategy is required to sample memory units from the memory bank. We adopt a greedy matching algorithm based on frame-pair similarity, where similarity is defined using the field-of-view (FOV) overlap ratio and timestamp differences as confidence measures. Algorithm 1 presents our approach to memory retrieval. Additionally, we apply a similarity filter to eliminate redundant units, ensuring that only the most relevant frames are retrieved for reference in the current generation. Figure 3 visualizes the overlapping metric with the Monte Carlo sampling.
Input:
  • A memory bank of $n$ historical states $\{(\mathbf{I}_{i},\mathbf{p}_{i},t_{i})\}_{i=1}^{n}$.
  • Current state $(\mathbf{I}_{c},\mathbf{p}_{c},t_{c})$.
  • Memory condition length $L$.
  • Similarity threshold $tr$.
  • Weights $w_{o},w_{t}$.
Output: A list of selected state indices $S$.
1
2begin
3       $\Delta$ Compute Confidence Score
4       Compute FOV overlap ratio $\mathbf{o}$ with the current state via Monte Carlo sampling.
5       Compute time difference $\mathbf{d}=\text{Concat}(\{|t_{i}-t_{c}|)\}_{i=1}^{n}$.
6       Compute confidence $\boldsymbol{\alpha}=\mathbf{o}\cdot w_{o}$ $\mathbf{d}\cdot w_{t}$.
7
8      $\Delta$ Selection with Similarity Filtering
9       Initialize $S=\varnothing$.
10       for $m=1$ to $L$ do
11             Select the state $i^{*}$ with the highest $\alpha_{i^{*}}$.
12             Append $i^{*}$ to $S$.
13             Remove all states $j$ whose feature similarity with $i^{*}$ exceeds $tr$.
14
15       end for
16      Return $S$.
17 end
18
@@ 8      $\Delta$ Selection with Similarity Filtering
Figure 3: Two-view FOV overlapping visualization
3.5 State Embedding Design
The design of state embedding is critical for the memory block to retrieve valid information effectively.
A key aspect of this process is capturing spatial information, which primarily depends on pose embeddings. Inspired by [14, 8], which transforms frame-wise poses into dense positional embeddings to encode fine-grained spatial details, we adopt Plücker embedding [30] as our pose representation. For timestamp embeddings, a simple MLP-based mapping is sufficient. The state embedding $\mathbf{E}$ is then computed as the sum of pose and timestamp embeddings:
$\mathbf{E}=G_{p}(\text{PE}(\mathbf{p}))+G_{t}(t),$ (8)
where $G_{p}$ and $G_{t}$ are MLP layers mapping vectors to embeddings, and PE denotes the Plücker embedding function, which maps $\mathbf{p}\in\mathbb{R}^{5}$ (x, y, z, pitch, yaw) to $\mathbf{p}\in\mathbb{R}^{h\times w\times 6}$.
Additionally, we observe that relative embeddings are more effective than absolute embeddings for spatial reasoning, as they significantly reduce the model’s learning complexity. To leverage this, we introduce relative embedding in the memory attention block, where the Query embedding is always set to zero, and the Key embedding is derived from the relative pose. To integrate relative embeddings into our architecture, we separate the cross-attention into frame-wise querying, where input frames attend to memory frames independently. This design is illustrated in Figure 2 (d).

### v2

@@ Abstract
World simulation has gained increasing popularity due to its ability to model virtual environments and predict the consequences of actions. However, the limited temporal context window often leads to failures in maintaining long-term consistency, particularly in preserving 3D spatial consistency. In this work, we present WorldMem, a framework that enhances scene generation with a memory bank consisting of memory units that store memory frames and states (e.g., poses and timestamps). By employing state-aware memory attention that effectively extracts relevant information from these memory frames based on their states, our method is capable of accurately reconstructing previously observed scenes, even under significant viewpoint or temporal gaps. Furthermore, by incorporating timestamps into the states, our framework not only models a static world but also captures its dynamic evolution over time, enabling both perception and interaction within the simulated world. Extensive experiments in both virtual and real scenarios validate the effectiveness of our approach. Project page at https://xizaoqu.github.io/worldmem.
Figure 1: WorldMem enables long-term consistent world generation with an integrated memory mechanism. (a) Previous world generation methods typically face the problem of inconsistent world due to limited temporal context window size. (b) WorldMem  empowers the agent to explore diverse and consistent worlds with an expansive action space, e.g., crafting environments by placing objects like pumpkin light or freely roaming around. Most importantly, after exploring for a while and glancing back, we find the objects we placed are still there, with the inspiring sight of the light melting the surrounding snow, testifying to the passage of time. Red and green boxes indicate scenes that should be consistent.
@@ 1 Introduction
World simulation has gained significant attention for its ability to model environments and predict the outcomes of actions (bar2024navigationworldmodels; oasis2024; alonso2025diffusion; feng2024matrix; parkerholder2024genie2; valevski2024diffusion). Recent advances in video diffusion models have further propelled this field, enabling high-fidelity rollouts of potential future scenarios based on user actions, such as navigating through an environment or interacting with objects. These capabilities make world simulators particularly promising for applications in autonomous navigation (feng2024matrix; bar2024navigationworldmodels) and as viable alternatives to traditional game engines (oasis2024; parkerholder2024genie2).
Despite these advances, a fundamental challenge remains: the limited probing horizon. Due to computational and memory constraints, video generative models operate within a fixed context window and are unable to condition on the full sequence of past generations. Consequently, most existing methods simply discard previously generated content, leading to a critical issue of world inconsistency, which is also revealed in wang2025error. As illustrated in Figure 1(a), when the camera moves away and returns, the regenerated content diverges from the earlier scene, violating the coherence expected in a consistent world.
A natural solution is to maintain an external memory that stores and retrieves relevant historical information outside the generative loop. While intuitive, formulating such a memory mechanism is non-trivial. A direct approach might involve explicit 3D scene reconstruction to preserve geometry and detail. However, 3D representations are inflexible in dynamic and evolving environments and are prone to loss of detail, especially for large, unbounded scenes (wu2025genfusion).
Instead, we argue that geometry-free representations offer a more flexible solution. These representations, however, pose their own challenges – particularly in balancing detail retention with memory scalability. For example, implicit approaches like storing abstract features via LoRA modules (hong2024slowfast) offer compactness but lose visual fidelity and spatial specificity. Some recent works represent visual scenes as discrete tokens encoding fine-grained visual information (sajjadi2022scene; jiang2025rayzer), but they are limited by a fixed token and struggle to capture the complexity of diverse and evolving environments. To address this issue, we observe that for generating the immediate future, only a small subset of historical content is typically relevant. Based on this, we propose a token-level memory bank that stores all previously generated latent tokens, and retrieves a targeted subset for each generation step based on relevance.
Conditioning on the retrieved memory requires spatial-temporal reasoning. In contrast to prior work where memory aids local temporal smoothness (zheng2024memo) or semantic coherence (wu2025corgi; rahman2023make), long-term world simulation demands reasoning over large spatiotemporal gaps, e.g., memory and query may differ in viewpoint and time, and retain exact scenes with detail. To facilitate this reasoning, we propose augmenting each memory unit with explicit state cues, including spatial location, viewpoint, and timestamp. These cues serve as anchors for reasoning and are embedded as part of the query-key attention mechanism. Through this state-aware attention, our model can effectively reason the current frame with past observations, facilitating accurate and coherent generation. Importantly, such a design leverages standard attention architectures, enabling it to scale naturally with modern hardware and model capacity.
Motivated by this idea, we build our approach, WorldMem, on top of the Conditional Diffusion Transformer (CDiT) (peebles2023scalable) and the Diffusion Forcing (DF) paradigm (chen2025diffusion), which autoregressively generates first-person viewpoints conditioned on external action signals. As discussed above, at the core of WorldMem is a memory mechanism composed of a memory bank and memory attention. To ensure efficient and relevant memory retrieval from the bank, we introduce a confidence-based selection strategy that scores memory units based on field-of-view (FOV) overlap and temporal proximity. In the memory attention, the latent tokens being generated act as queries, attending to the memory tokens (as keys and values) to incorporate relevant historical context. To ensure robust correspondence across varying viewpoints and time gaps, we enrich both queries and keys with state-aware embeddings. A relative embedding design is introduced to ease the learning of spatial and temporal relationships. This pipeline enables precise, scalable reasoning over long-range memory, ensuring consistency in dynamic and evolving world simulations.
@@
We evaluate WorldMem on a customized Minecraft benchmark (fan2022minedojo) and on RealEstate10K (zhou2018stereo). The Minecraft benchmark includes diverse terrains (e.g., plains, savannas, and deserts) and various action modalities (movement, viewpoint control, and event triggers), which is a wonderful environment for idea verification. Extensive experiments show that WorldMem significantly improves 3D spatial consistency, enabling robust viewpoint reasoning and high-fidelity scene generation, as shown in Figure 1(b). Furthermore, in dynamic environments, WorldMem accurately tracks and follows evolving events and environment changes, demonstrating its ability to both perceive and interact with the generated world. We hope our promising results and scalable designs will inspire future research on memory-based world simulation.
@@ 2 Related Work
Video diffusion model. With the rapid advancement of diffusion models (song2020score; peebles2023scalable; chen2025diffusion), video generation has made significant strides (wang2023modelscope; wang2023lavie; chen2023videocrafter1; guo2023animatediff; sora; jin2024pyramidal; yin2024slow). The field has evolved from traditional U-Net-based architectures (wang2023modelscope; chen2023videocrafter1; guo2023animatediff) to Transformer-based frameworks (sora; ma2024latte; opensora), enabling video diffusion models to generate highly realistic and temporally coherent videos. Recently, autoregressive video generation (chen2025diffusion; kim2024fifodiffusion; henschel2024streamingt2v) has emerged as a promising approach to extend video length, theoretically indefinitely. Notably, Diffusion Forcing (chen2025diffusion) introduces a per-frame noise-level denoising paradigm. Unlike the full-sequence paradigm, which applies a uniform noise level across all frames, per-frame noise-level denoising offers a more flexible approach, enabling autoregressive generation.
@@
Interactive world simulation. World simulation aims to model an environment by predicting the next state given the current state and action. This concept has been extensively explored in the construction of world models (ha2018world) for agent learning (ha2018recurrent; hafner2019dream; hafner2020mastering; hu2023gaia; beattie2016deepmind; yang2023learning). With advances in video generation, high-quality world simulation with robust control has become feasible, leading to numerous works focusing on interactive world simulation (bar2024navigationworldmodels; oasis2024; alonso2025diffusion; feng2024matrix; parkerholder2024genie2; valevski2024diffusion; yu2025gamefactory; yu2025survey; yu2025position). These approaches enable agents to navigate generated environments and interact with them based on external commands.
@@
However, due to context window limitations, such methods discard previously generated content, leading to inconsistencies in the simulated world, particularly in maintaining 3D spatial coherence.
@@
Consistent world simulation. Ensuring the consistency of a generated world is crucial for effective world simulation wang2025error. Existing approaches can be broadly categorized into two types: geometric-based and geometric-free. The geometric-based methods explicitly reconstruct the generated world into a 3D/4D representation (liu2024reconx; gao2024cat3d; wang20243d; ren2025gen3c; yu2024wonderjourney; yu2024wonderworld; liang2024wonderland). While this strategy can reliably maintain consistency, it imposes strict constraints on flexibility: Once the world is reconstructed, modifying or interacting with it becomes challenging. Geometric-free methods focus on implicit learning. Methods like alonso2025diffusion; valevski2024diffusion ensure consistency by overfitting to predefined scenarios (e.g., specific CS:GO or DOOM maps), limiting scalability. StreamingT2V (henschel2024streamingt2v) maintains long-term consistency by continuing on both global and local visual contexts from previous frames, while SlowFastGen (hong2024slowfast) progressively trains LoRA (hu2022lora) modules for memory recall. However, these methods rely on abstract representations, making accurate scene reconstruction challenging. In contrast, our approach retrieves information from previously generated frames and their states, ensuring world consistency without overfitting to specific scenarios.
@@ 3 WorldMem
This section details the methodology of WorldMem. Sec. 3.1 introduces the relevant preliminaries, while Sec. 3.2 describes the interactive world simulator serving as our baseline. Sec. 3.3 and  3.4 present the core of our proposed memory mechanism.
@@ 3.1 Preliminary
Autoregressive video generation. Autoregressive video generation aims to extend videos over the long term by predicting frames sequentially (videopoet; wu2023ardiffusion). While various methods exist for autoregressive generation, Diffusion Forcing (DF) (chen2025diffusion) provides a neat and effective approach to achieve this. Specifically, DF introduces per-frame noise levels $k_{t}$:
@@ 3.2 Interactive World Simulation
Before introducing the memory mechanism, we first present our interactive world simulator, which models long video sequences using an auto-regressive conditional diffusion transformer. Interaction is achieved by embedding external control signals, primarily actions, into the model through dedicated conditioning modules (parkerholder2024genie2; oasis2024; yu2025gamefactory).
@@
Following prior work (oasis2024), we adopt a conditional Diffusion Transformer (DiT) (peebles2023scalable) architecture for video generation, and Diffusion Forecasting (DF) (chen2025diffusion) for autoregressive prediction. As shown in Figure 2(a), our model consists of multiple DiT blocks with spatial and temporal modules for spatiotemporal reasoning. The temporal module applies causal attention to ensure that each frame only attends to preceding frames.
The actions are injected by first projected into the embedding space using a multi-layer perceptron (MLP). The resulting action embeddings are added to the denoising timestep embeddings and injected into the temporal blocks using Adaptive Layer Normalization (AdaLN) (xu2019understanding), following the paradigm of bar2024navigationworldmodels; oasis2024. In our Minecraft experiments, the action space contains 25 dimensions, including movements, view adjustments, and event triggers. We also apply timestep embeddings to the spatial blocks in the same manner, although this is omitted from the figure for clarity. Standard architectural components such as residual connections, multi-head attention, and feedforward networks are also not shown.
@@
The combination of conditional DiT and DF provides a strong baseline for long-term interactive video generation. However, due to the computational cost of video synthesis, the temporal context window remains limited. As a result, content outside this window is forgotten, which leads to inconsistencies during long-term generation (oasis2024).
Input: Memory bank of $N$ historical states $\{(\mathbf{x}^{m}_{i},\mathbf{p}_{i},t_{i})\}_{i=1}^{N}$;
Current state $(\mathbf{x}_{c},\mathbf{p}_{c},t_{c})$; memory condition length $L_{M}$;
Similarity threshold trtr; weights $w_{o}$, $w_{t}$.
Output: A list of selected state indices $S$
Compute Confidence Score:
Compute FOV overlap ratio $\mathbf{o}$ via Monte Carlo sampling.
Compute time difference $\mathbf{d}=\text{Concat}(\{|t_{i}-t_{c}|\}_{i=1}^{n})$.
Compute confidence $\boldsymbol{\alpha}=\mathbf{o}\cdot w_{o}-\mathbf{d}\cdot w_{t}$.
[4pt]
Selection with Similarity Filtering:
Initialize $S=\varnothing$
for $m=1$ to $L_{M}$ do
      Select $i^{*}$ with highest $\alpha_{i^{*}}$
      Append $i^{*}$ to $S$
[…]
return $S$
@@
3.3 Memory Representation and Retrieval
To address the limited context window of video generative models, we introduce a memory mechanism that enables the model to retain and retrieve information beyond the current generation window. This mechanism maintains a memory bank composed of historical frames and their associated state information: $\{(\mathbf{x}^{m}_{i},\mathbf{p}_{i},t_{i})\}_{i=1}^{N},$ where $\mathbf{x}^{m}_{i}$ denotes a memory frame, $\mathbf{p}_{i}\in\mathbb{R}^{5}$ (x, y, z, pitch, yaw) is its pose, and $t_{i}$ is the timestamp. Each tuple is referred to as a memory unit. We save $\mathbf{m}_{i}$ in token-level, which is compressed by the visual encoder but retains enough details for reconstruction. The corresponding states $\{(\mathbf{p},t)\}$ play a critical role not only in memory retrieval but also in enabling state-aware memory conditioning.
Memory Retrieval. Since the number of memory frames available for conditioning is limited, an efficient strategy is required to sample memory units from the memory bank. We adopt a greedy matching algorithm based on frame-pair similarity, where similarity is defined using the field-of-view (FOV) overlap ratio and timestamp differences as confidence measures. Algorithm 1 presents our approach to memory retrieval. Although simple, this strategy proves effective in retrieving relevant information for conditioning. Moreover, the model’s reasoning over memory helps maintain performance even when the retrieved content is imperfect.
3.4 State-aware Memory Condition
After retrieving necessary memory units, unlike prior methods that use memory mainly for temporal smoothness (zheng2024memo) or semantic guidance (wu2025corgi; rahman2023make), our goal is to explicitly reconstruct previously seen visual content – even under significant viewpoint or scene changes. This requires the model to perform spatiotemporal reasoning to extract relevant information from memory, which we model using cross-attention (vaswani2017attention). Since relying solely on visual tokens can be ambiguous, we incorporate the corresponding states as cues to enable state-aware attention.
State Embedding. State embedding provides essential spatial and temporal context for memory retrieval. To encode spatial information, we adopt Plücker embedding (sitzmann2021light) to convert 5D poses $\mathbf{p}\in\mathbb{R}^{5}$ into dense positional features $\text{PE}(\mathbf{p})\in\mathbb{R}^{h\times w\times 6}$, following (he2024cameractrl; gao2024cat3d). Temporal context is captured via a lightweight MLP over sinusoidal embedded (SE) timestamps. The final embedding is (Figure 2 (c)):
$\mathbf{E}=G_{p}(\text{PE}(\mathbf{p}))+G_{t}(\text{SE}(t)),$ (3)
where $G_{p}$ and $G_{t}$ are MLPs mapping pose and time into a shared space.
State-aware Memory Attention. To support reconstruction under viewpoint and temporal shifts, we introduce a state-aware attention mechanism that incorporates spatial-temporal cues into memory retrieval. By conditioning attention on both visual features and state information, the model achieves more accurate reasoning between input and memory.
Let $\mathbf{X}_{q}\in\mathbb{R}^{l_{q}\times d}$ denote the flattened feature map of input frames (queries), and $\mathbf{X}_{k}\in\mathbb{R}^{l_{k}\times d}$ the concatenated memory features (keys and values). We first enrich both with their corresponding state embeddings $\mathbf{E}_{q}$ and $\mathbf{E}_{k}$:
$\tilde{\mathbf{X}}_{q}=\mathbf{X}_{q}+\mathbf{E}_{q},\quad\tilde{\mathbf{X}}_{k}=\mathbf{X}_{k}+\mathbf{E}_{k}.$ (4)
Cross-attention is then applied to retrieve relevant memory content and output updated $\mathbf{X}^{\prime}$:
$\mathbf{X}^{\prime}=\text{CrossAttn}(Q=p_{q}(\tilde{\mathbf{X}}_{q}),\ K=p_{k}(\tilde{\mathbf{X}}_{k}),V=p_{v}(\mathbf{X}_{k})),$ (5)
where $p_{q}$, $p_{k}$, and $p_{v}$ are learnable projections.
To simplify the reasoning space, we adopt a relative state formulation. For each query frame, the state is set to a zero reference (e.g., the pose is reset to the identity and the timestamp to zero), while the states of key frames are normalized to relative values. This design, illustrated in Figure 2(d), improves alignment under viewpoint changes and simplifies the learning objective.
Incorporating memory into pipeline. We incorporate memory frames into the pipeline by treating them as clean inputs during both training and inference. As shown in Figure 2 (a-b), during training, memory frames are assigned the lowest noise level $k_{\text{min}}$, while context window frames receive independently sampled noise levels from the range $k_{\text{min}},k_{\text{max}}]$. During inference, both memory and context frames are assigned $k_{\text{min}}$, while the current generating frames are assigned $k_{\text{max}}$.
To restrict memory influence only to memory blocks, we apply a temporal attention mask:
$A_{\text{mask}}(i,j)=\begin{cases}1,&i\leq L_{M}\ \text{and}\ j=i\\ 1,&i>L_{M}\ \text{and}\ j\leq i\\ 0,&\text{otherwise}\end{cases}$ (6)
where $L_{M}$ is the number of memory frames that are appended before frames within the context window. This guarantees causal attention while preventing memory units from affecting each other.